OpenStudy (anonymous):

subtract: (3/x-3) - (5/x-2)

OpenStudy (campbell_st):

you need a common denominator of (x -3)(x - 2) so you have \[\frac{3(x -2)}{(x -3)(x-2)} - \frac{5(x -3)}{(x -3)(x-2)}\] just distribute the numerators and collect like terms... then you'll have an answer

OpenStudy (anonymous):

thnks so much

OpenStudy (campbell_st):

so it can be written as \[\frac{3(x -2) - 5(x -3)}{(x -3)(x-2)}\]

OpenStudy (anonymous):

so.. 2x+9/(x-3)(X-2)

OpenStudy (phi):

when you distribute the 3 and the -5 the top becomes 3x -6 -5x +15 how do you get +2x ? Isn't it -2x +9

OpenStudy (anonymous):

i didnt put a +2

OpenStudy (phi):

what is this? so.. 2x+9/(x-3)(X-2)

OpenStudy (anonymous):

ya thts a plus 9

OpenStudy (anonymous):

nevermind